OpenStudy (anonymous):

There are 23 coins in a bank. All the coins are dimes and quarters. The total value of the coins is $3.80. How many dimes are there? How many quarters? Write the system of equations that would be used to solve this problem.

OpenStudy (anonymous):

4x = 10y 4x + 10y = 3.8 x + y = 23 now just solve it, :))

OpenStudy (anonymous):

yaay

OpenStudy (valpey):

10d+25q = 380 d+q=23 -10d-10q=-230 10d-10d+25q-10q = 380-230 15q = 150 q = 10 d = 13